A CAR IS TRAVELING WITH A LINEAR VELOCITY v ON A CIRCULAR ROAD OF RADIUS R IF IT IS INCREASING ITS SPEED AT THE RATE OF 'a' MS^-2 THEN THE RESULTANT ACCELERATION WILL BE

Grade:12

1 Answers

Ramesh V
14 years ago

 

The instantaneous acceleration ar pointing towards the center of the circle is called radial acceleration or centripetal acceleration.                      The radial component causes a change in direction and its magnitude is given by ar = v2/r, where r is the radius of curvature at the point in question.

The tangential component causes a change in speed and its magnitude is given by at = dv/dt = a m/sec2 (given)

The tangential acceleration of the car has a constant magnitude of at = a m/sec2

The components ar and at are perpendicular to each other and the magnitude of a is given by

resultant accln. a2 = ar2 + at2

                         resultant accln. (ares)         =[ a2 +(v2/r)2 ]1/2
                                                                       =[ a2 +(v4/r2) ]1/2

---

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch